can you help me with this problem.

If 2(x+3) = x + 10, the x equals
thank you
julie

Answer by checkley71(8403)

2(x+3)=x+10
2x+6=x+10
2x-x=10-6
x=4 answer.
proof
2(4+3)=4+10
2*7=14
14=14